OpenStudy (anonymous):

In the equation =(x+h)^2 How does that distribute out to =x^2+2xh+h^2 I guess my main question is where does the 2xh come from?? Thank You

OpenStudy (perl):

(x+h)(x+h) = x*x + h*x + x*h + h*h = x^2 + 2hx + h^2

OpenStudy (some.random.cool.kid):

x is the varible it is also what is known as the x axis can be used as explained @above